OpenStudy (anonymous):

The length of a rectangle is 8 cm greater than it's width. Find the dimensions of the rectangle if its area is 150 cm^2.

OpenStudy (anonymous):

Hmm are you supposed to solve or make a system of equation? (:

OpenStudy (amistre64):

a(a+8) = 150 a^2 +8a = 150 (a+4)^2 = 166 a = -4 +- sqrt(166)

OpenStudy (amistre64):

thats just odd lol

OpenStudy (amistre64):

150 / 16.8841 = 8.88409806